OpenStudy (anonymous):

Let f(x) = x^2, evaluate lim f(h-2)-f(-2) ---- ---------- h->0 h

OpenStudy (anonymous):

How far did you get?

OpenStudy (anonymous):

I can't figure it out.

OpenStudy (anonymous):

Did you start?

OpenStudy (anonymous):

No? I don't understand it.

OpenStudy (anonymous):

if \[f(x) = x^2\] What is f(h-2) ? What is f(-2) ? What is \[f(h-2) - f(-2) \over h \] ?

OpenStudy (anonymous):

This is basic algebra type stuff. The 'I don't get it' response doesn't quite cut it. What don't you get?

OpenStudy (anonymous):

I just need help....I don't need a lecture. Thank you for attempting to help, but I'd rather not have to further explain myself and beg for help from someone who does not want to give it.

OpenStudy (anonymous):

I have 174 fans and 218 medals. I give help. What I don't give are answers. If you want to learn, answer my questions. If you want answers then be prepared to struggle when it comes time to take the test.

OpenStudy (anonymous):

I don't want answers, I want someone to explain it? Whatever. Please leave me question page. Thank you.

OpenStudy (anonymous):

I did explain it. If f(x) = x^2 what is f(-2) ?

OpenStudy (anonymous):

4............but then you were rude about it.

OpenStudy (anonymous):

And what is f(h-2) ?

OpenStudy (anonymous):

h-4? See I don't UNDERSTAND.

OpenStudy (anonymous):

How did you find f(-2) ?

OpenStudy (anonymous):

I don't know.............

OpenStudy (anonymous):

Please just leave my page.

OpenStudy (anonymous):

You got the answer. You just lack confidence. It isn't hard. f(-2) = (-2)^2, f(a) = a^2, f(h-2) = (h-2)^2... Engage your brain a bit and think about what you know.

OpenStudy (anonymous):

Or not. It's your grade I guess.

OpenStudy (anonymous):

Leave please.